graph the equation 6x+6y=30 by finding the intercepts.
My teacher has given us the answers, but I do not understand how I am supposed to figure this out without just copying the answer.

Found 2 solutions by rfer, mananth:
Answer by rfer(16322)

x intercept is when y=0
y intercept is when x=0
------------
6x+6Y=30
6x+6*0=30
6x+0=30
x=30/6
x=5
(5,0)
------------
because the numbers are the same
y intercept = (0,5)

Answer by mananth(16949)

6x+6y=30
divide equation by 6
x+y=5
y intercept point
when x=0, y=5 (0,5)
X intercept point
when y = 0 , x=5, (5,0)
Plot the points draw the line graph

m.ananth@hotmail.ca